# Knowledge Distillation with Structured Chain-of-Thought for Text-to-SQL

### v1

@@ 3.2 Analysis of Model Failures
The unstructured distillation (ReasonSQL) baseline evaluates distillation through an unstructured reasoning trace (). As shown in Figure 3(d), this approach achieved an execution accuracy of , higher than that of and FN-Gold, indicating that unstructured CoT reasoning provides a better distillation signal than both the naive model and FN-Gold. ReasonSQL significantly reduced the rate of generation errors compared to the naive Student Model , suggesting that unstructured distillation enhanced the stability of the model. Furthermore, it reduced semantic errors from 53.8% to 39.8% (a -point reduction) and reduced syntactic errors from to (a -point reduction), primarily by improving basic schema linking and eliminating ‘No Such Table’ errors. These findings align with previous research demonstrating that CoT-based distillation improves accuracy [hsieh2023cotdistilling] and further clarify its specific advantages.
@@ 3.3 Fine-Grained Performance Analysis and Ablation Study
To investigate the precise source of performance of Struct-SQL, we move beyond aggregate metrics to examine the detailed differences between models.
Performance by SQL Construct: Assessing model performance across the spectrum of query formulations can be insightful; therefore, we analyzed execution accuracy for key SQL constructs, as shown in Figure 4(a). The comparative analysis indicates that Struct-SQL’s performance gains are most pronounced in tasks that require aggregation and explicit structural decomposition. This is evidenced by its strong results on specific SQL operations. For queries requiring a GROUP BY clause, the Struct-SQL model ( EX) outperformed both FN-Gold ( EX) and ReasonSQL ( EX), demonstrating that explicitly distilling the query execution plan effectively trains the model on the necessary aggregation formulation. An advantage was also observed for queries requiring an ORDER BY, where Struct-SQL showed a -point absolute improvement over ReasonSQL. However, the comparative analysis revealed a notable exception: for queries that require JOIN operations, ReasonSQL achieved the highest execution accuracy. In contrast, Struct-SQL’s score ( EX) was lower, suggesting that the unstructured CoT trace provided a unique advantage for learning multi-entity linking. This area remains a core challenge for all models, as confirmed by the Teacher Model’s performance ( EX). Furthermore, all baselines, including Struct-SQL, shared a weakness in handling queries that require Set Operations.
@@
Evaluation of Knowledge Transfer: Gains vs. Losses: A Gains vs. Losses analysis was conducted to assess the quality of the distilled knowledge (Figure 4(b)). This approach compares queries where the Teacher was correct, but the Student Model failed (Losses, red segments) with cases where the Student Model was correct, but the Teacher Model was incorrect (Gains, green segments), representing true generalization. The results indicate that Struct-SQL demonstrates the highest net performance relative to the Teacher Model by exhibiting the fewest losses and the most gains among all student models. This confirms that the structured reasoning signal leads to better fidelity in knowledge transfer.
@@
Performance State Transitions: Figure 4(c) illustrates the progression of queries from the baseline Student Model to the distilled Struct-SQL, highlighting the systematic conversion of error states into successful or unsuccessful output. Struct-SQL retained more than 81% of the initial successes and enhanced stability by reducing generation errors of the Student Model from 4.0% to 0.4%. The structured framework also corrected 41.3% of semantic errors (SEM SUCC), confirming that the formal query plan addresses logical flaws. Additionally, over two-thirds of syntactic failures were upgraded: 29% were corrected (SYN SUCC) and 41% shifted to semantic errors (SYN SEM).
@@
Generalization on Mistral-7B: To validate the transferability of our framework to other architectures, we replicated our experiments on Mistral-7B-Instruct-v3.0. This model was specifically selected for its low zero-shot accuracy of 4.4% EX on BIRD mini-dev. As shown in Table 4, Struct-SQL shows generalizability, achieving a performance advantage (29.3% EX) over the Student Model (7.2% EX) and the ReasonSQL baseline (25.1% EX), confirming that the structured query plan offers a robust supervision signal independent of the base model.
@@
Computational Efficiency: On a single H200 GPU, using Qwen3-4B-Instruct-2507 as the Student Model, Struct-SQL converged using an early stopping strategy with a patience of 8 steps and a threshold of 0.001 in 2.24 epochs, requiring only 29.15 minutes. This training efficiency is comparable to that of the unstructured ReasonSQL, which required 25.24 minutes across 6.4 epochs with the same batch size, sample count and early stopping strategy. This 1,000-sample distillation strategy is also computationally tractable compared to traditional finetuning on larger datasets. For context, the FN-Gold baseline, trained on the entire 9,000+ sample BIRD dataset, required 110.57 minutes (4.33 epochs) to converge. This efficiency underscores the practical deployability of our structured distillation method in resource-constrained environments.
@@
Official Test Evaluation: Verified by the BIRD benchmark team, our model achieves 60.42% EX333BIRD public leaderboard listing is deferred to maintain double-blind anonymity., ranking first among B models (as of December 18, 2025). This result is achieved using strict single-model inference with greedy decoding and no self-consistency.
@@ 4.1 Text-to-SQL with In-Context Learning (ICL)
LLMs fundamentally redefined the field of Text-to-SQL [deng2022recent], achieving substantial performance on challenging cross-domain benchmarks such as Spider [yu2018spider] and BIRD [li2024can]. ICL [dong-etal-2024-survey] has emerged as a dominant research paradigm, enabling LLMs to leverage their pre-training by processing instructions and examples for SQL generation [shi2025asurvey]. Initially, ICL approaches utilized foundational zero-shot and few-shot paradigms [dong2023zero, rajkumar2022evaluating, liu2023comprehensive, wang-etal-2025-mac]. However, generating complex SQL queries that involve nested logic and multitable joins was found to require explicit external guidance [pourreza2023dinsql]. As a result, the field pivoted toward decomposition, a multi-step reasoning tactic designed to improve accuracy and reliability [pourreza2023dinsql, zhang2023act, gao2024dail, Li2024pet, liu2023divide]. As a form of ICL-based decomposition, basic CoT helped LLMs break down the task, follow a logical workflow and improve execution accuracy [liu2023divide]. This progression led to sophisticated variants such as DIN-SQL [pourreza2023dinsql], which decomposes the task into sequential stages (e.g., schema linking, query classification). Additional multi-stage architectural interventions have also been explored [zhang2023act, Li2024pet]. Crucially, these approaches often require multiple LLM calls to intervene, correct, or gather intermediate data, leading to high latency and cost. To mitigate this, some approaches focused on optimizing efficiency by creating single-pass Structural CoT approaches [pourreza2024chase, liu2023divide]. These methods instruct the LLM to follow a formalized logical blueprint to gain the necessary understanding of the schema and natural language before generating the SQL. Specialized methods, including QP-CoT and QDecompose, excel here, as they have been shown to produce better results than non-structured CoT [pourreza2024chase, gao2024dail]. For example, QP-CoT provides few-shot query plan examples that require the model to first generate a detailed query plan before generating SQL code [pourreza2024chase]. Although single-pass ICL performs well with large LLMs, its gains are limited when applied to SLMs [maamari2024death] or require expensive finetuning [pourreza2024dts, liu2025uncovering, gorti-etal-2025-msc, zhai2025excot]. This work addresses this identified gap by introducing Struct-SQL, a novel KD method that enables single-pass ICL for SLMs.
@@ 4.2 KD for Reasoning Transfer
KD has evolved from a model compression technique to a method for complex knowledge transfer, called Skill Distillation, often used to transfer reasoning knowledge from one model to others or to facilitate self-improvement [hinton2014distilling, shi2025asurvey]. Due to the significant resource demands of LLMs, recent research has focused on efficiently transferring its reasoning capabilities from a teacher LLM to a student SLM. This transfer is frequently grounded in ICL principles, demonstrating that learning from a few-shot demonstrations can be successfully distilled into the SLMs’ parameters [snell2022learning]. The primary established approach to enable reasoning in SLMs is the Finetune-CoT paradigm [li2024explanations], which uses unstructured CoT explanations of the teacher as a supervision signal. This distillation approach has been demonstrated to be effective with multi-step mathematical reasoning [yao2023specializing]. However, relying on unstructured CoT often leads SLMs to learn spurious correlations rather than deep causal features, thereby reducing robustness on complex data [li2023mixed, shridhar2023distilling]. Consequently, recent research increasingly favors methods that enforce structural decomposition and explicit causality. These approaches include creating constrained distillation pipelines, with the aim of removing ambiguous input context through advanced structural formats. For example, SocraticCoT provides an explicit structural format by guiding the model through reasoning using defined subquestion-solution pairs [shridhar2023distilling]. Alternatively, Mixed Distillation integrates the CoT with a formal verifiable Program-of-Thought reasoning to provide a less ambiguous supervision signal [li2023mixed]. In the Text-to-SQL domain, structured KD has seen limited uptake. Previous studies, while validating the effectiveness of distillation in this area, have generally relied on less structured signals, such as schema-based finetuning to improve schema linking [hong2024next], the inclusion of enterprise-specific custom examples [hoang-etal-2025-distill], and unstructured CoT reasoning transfer [rossiello2025rationalization]. The proposed Struct-SQL framework systematically evaluates whether structure-based reasoning transfer, specifically QP-CoT, improves the distillation for Text-to-SQL.
@@ 5 Conclusion and Future Research
This work presents Struct-SQL, a KD framework that transfers structured reasoning (QP-CoT) from a Teacher LLM to a smaller student model for the Text-to-SQL task. The central hypothesis is that a formal, structured reasoning signal provides a superior, less ambiguous teaching blueprint than unstructured reasoning. The experiments confirm the hypothesis, demonstrating that distilling structured reasoning is a better teaching method. Detailed error analysis indicates that these improvements are primarily due to a significant reduction in syntactic errors. An additional ablation study demonstrated generalization to a different base model. This increased precision, however, comes with a trade-off: the mandatory generation of the intermediate query plan increases the inference token overhead and relies on a fixed reasoning template. Additionally, the Student’s performance remains bound by the Teacher’s fidelity, particularly with respect to complex join and set operations. Future research should therefore explore integrating or combining this structured reasoning with other established strategies to determine the optimal signal for distillation efficiency. Given the demonstrated effectiveness of structured signals in the refinement of syntactic efficiency and logical reasoning, further research is warranted to validate the effect of this structured distillation approach on other complex reasoning tasks beyond Text-to-SQL. In summary, the proposed framework facilitates the deployment of high-performing, cost-effective and private models.

### v2

@@ 3.1 Overall Performance
| ReasonSQL (mismatched)222Distilled using CoT, evaluated with QP-CoT. | QP-CoT | 29.20% | 46.62% | 24.80% | 14.71% | |
@@
To isolate whether Struct-SQL’s performance gains stem from the structured QP-CoT prompt format at inference time or from the structured supervisory signal during training, we conducted an additional ablation study. We evaluated the ReasonSQL model—trained on unstructured CoT traces—using the QP-CoT prompt at inference, denoted as ReasonSQL (mismatched) in Table 3. This configuration achieved only 29.2% execution accuracy, a substantial drop from the 36.9% achieved when ReasonSQL uses its native CoT prompt. This 7.7-point degradation reveals a critical prompt-training mismatch: a model trained on unstructured reasoning cannot effectively leverage a structured prompt format, even when the prompt includes few-shot examples demonstrating how to generate the structured query plan. This suggests that in SLMs, ICL alone is insufficient to bridge the gap; rather, the model must be explicitly trained on structured reasoning to internalize the required logical decomposition patterns. The strength of the training signal is further evidenced by Struct-SQL, which achieves 45.0% EX using the identical QP-CoT prompt. This disparity validates that Struct-SQL’s success stems from internalizing the logical decomposition patterns during training, which the prompt alone cannot replicate.
@@ 3.2 Analysis of Model Failures
The unstructured distillation (ReasonSQL) baseline evaluates distillation through an unstructured reasoning trace (). As shown in Figure 3(d), this approach achieved an execution accuracy of , higher than that of and FN-Gold, indicating that unstructured CoT reasoning provides a better distillation signal than both the naive model and FN-Gold. ReasonSQL significantly reduced the rate of generation errors compared to the naive Student Model , suggesting that unstructured distillation enhanced the stability of the model. Furthermore, it reduced semantic errors from 53.8% to 39.8% (a -point reduction) and reduced syntactic errors from to (a -point reduction), primarily by improving basic schema linking and eliminating ‘No Such Table’ errors. These findings align with previous research demonstrating that CoT-based distillation improves accuracy [12] and further clarify its specific advantages.
@@ 3.3 Fine-Grained Performance Analysis and Ablation Study
To investigate the precise source of performance of Struct-SQL, we move beyond aggregate metrics to examine the detailed differences between models. In this section, we deconstruct the model’s capabilities across three dimensions: its proficiency with complex SQL operations (Figure 4(a)), the fidelity of its alignment with the teacher model (Figure 4(b)), and its ability to systematically rectify the student model’s baseline errors (Figure 4(c)).
3.3.1 Performance by SQL Construct:
Assessing model performance across the spectrum of query formulations can be insightful; therefore, we analyzed execution accuracy for key SQL constructs, as shown in Figure 4(a). The comparative analysis indicates that Struct-SQL’s performance gains are most pronounced in tasks that require aggregation and explicit structural decomposition. This is evidenced by its strong results on specific SQL operations. For queries requiring a GROUP BY clause, the Struct-SQL model ( EX) outperformed both FN-Gold ( EX) and ReasonSQL ( EX), demonstrating that explicitly distilling the query execution plan effectively trains the model on the necessary aggregation formulation. An advantage was also observed for queries requiring an ORDER BY, where Struct-SQL showed a -point absolute improvement over ReasonSQL. However, the comparative analysis revealed a notable exception: for queries that require JOIN operations, ReasonSQL achieved the highest execution accuracy. In contrast, Struct-SQL’s score ( EX) was lower, suggesting that the unstructured CoT trace provided a unique advantage for learning multi-entity linking. This area remains a core challenge for all models, as confirmed by the Teacher Model’s performance ( EX). Furthermore, all baselines, including Struct-SQL, shared a weakness in handling queries that require Set Operations.
@@ 3.3.1 Performance by SQL Construct:
A Gains vs. Losses analysis was conducted to assess the quality of the distilled knowledge (Figure 4(b)). This approach compares queries where the Teacher was correct, but the Student Model failed (Losses, red segments) with cases where the Student Model was correct, but the Teacher Model was incorrect (Gains, green segments), representing true generalization. The results indicate that Struct-SQL demonstrates the highest net performance relative to the Teacher Model by exhibiting the fewest losses and the most gains among all student models. This confirms that the structured reasoning signal leads to better fidelity in knowledge transfer.
3.3.3 Performance State Transitions:
@@
Figure 4(c) illustrates the progression of queries from the baseline Student Model to the distilled Struct-SQL, highlighting the systematic conversion of error states into successful or unsuccessful output. Struct-SQL retained more than 81% of the initial successes and enhanced stability by reducing generation errors of the Student Model from 4.0% to 0.4%. The structured framework also corrected 41.3% of semantic errors (SEM SUCC), confirming that the formal query plan addresses logical flaws. Additionally, over two-thirds of syntactic failures were upgraded: 29% were corrected (SYN SUCC) and 41% shifted to semantic errors (SYN SEM).
3.4 Generalization on Mistral-7B:
@@
To validate the transferability of our framework to other architectures, we replicated our experiments on Mistral-7B-Instruct-v3.0. This model was specifically selected for its low zero-shot accuracy of 7.2% EX on BIRD mini-dev with QP-CoT prompt. As shown in Table 4, Struct-SQL shows generalizability, achieving a performance advantage (29.3% EX) over the Student Model (7.2% EX) and the ReasonSQL baseline (25.1% EX), confirming that the structured query plan offers a robust supervision signal independent of the base model.
3.5 Computational Efficiency:
@@
On a single H200 GPU, using Qwen3-4B-Instruct-2507 as the Student Model, Struct-SQL converged using an early stopping strategy with a patience of 8 steps and a threshold of 0.001 in 2.24 epochs, requiring only 29.15 minutes. This training efficiency is comparable to that of the unstructured ReasonSQL, which required 25.24 minutes across 6.4 epochs with the same batch size, sample count and early stopping strategy. This 1,000-sample distillation strategy is also computationally tractable compared to traditional finetuning on larger datasets. For context, the FN-Gold baseline, trained on the entire 9,000+ sample BIRD dataset, required 110.57 minutes (4.33 epochs) to converge. This efficiency underscores the practical deployability of our structured distillation method in resource-constrained environments.
3.6 Official Test Evaluation:
@@
As shown in Table 4, Struct-SQL obtained 60.42% EX on the non-public BIRD benchmark [15], securing the top position among 4B models (as of January 30, 2026) in single model inference. This result is achieved using strict single-model inference with greedy decoding and no self-consistency.
| Model | Simple (%) | Moderate (%) | Challenging (%) | Overall EX (%) |
| Struct-SQL(4B) | 69.02 | 54.41 | 43.51 | 60.42 |
@@ 4.1 Text-to-SQL with In-Context Learning (ICL)
LLMs fundamentally redefined the field of Text-to-SQL [1], achieving substantial performance on challenging cross-domain benchmarks such as Spider [35] and BIRD [15]. ICL [3] has emerged as a dominant research paradigm, enabling LLMs to leverage their pre-training by processing instructions and examples for SQL generation [28]. Initially, ICL approaches utilized foundational zero-shot and few-shot paradigms [4, 26, 18, 31]. However, generating complex SQL queries that involve nested logic and multitable joins was found to require explicit external guidance [23]. As a result, the field pivoted toward decomposition, a multi-step reasoning tactic designed to improve accuracy and reliability [23, 37, 6, 17, 20]. As a form of ICL-based decomposition, basic CoT helped LLMs break down the task, follow a logical workflow, and improve execution accuracy [20]. This progression led to sophisticated variants such as DIN-SQL [23], which decomposes the task into sequential stages (e.g., schema linking, query classification). Additional multi-stage architectural interventions have also been explored [37, 17]. Crucially, these approaches often require multiple LLM calls to intervene, correct, or gather intermediate data, leading to high latency and cost. To mitigate this, some approaches focus on optimizing efficiency by creating single-pass Structural CoT approaches [22, 20]. These methods instruct the LLM to follow a formalized logical blueprint to gain the necessary understanding of the schema and natural language before generating the SQL. Specialized methods, including QP-CoT and QDecompose, excel here, as they have been shown to produce better results than non-structured CoT [22, 6]. For example, QP-CoT provides few-shot query plan examples that require the model to first generate a detailed query plan before generating SQL code [22]. Although single-pass ICL performs well with large LLMs, its gains are limited when applied to SLMs [21] or require expensive finetuning [24, 19, 7, 36]. This work addresses this identified gap by introducing Struct-SQL, a novel KD method that enables single-pass ICL for SLMs.
@@ 4.2 KD for Reasoning Transfer
KD has evolved from a model compression technique to a method for complex knowledge transfer, called Skill Distillation, which is often used to transfer reasoning knowledge from one model to another or to facilitate self-improvement [9, 28]. Due to the significant resource demands of LLM, recent research has focused on efficiently transferring their reasoning capabilities from a teacher LLM to a student SLM. This transfer is frequently grounded in ICL principles, demonstrating that learning from few-shot demonstrations can be successfully distilled into the SLMs’ parameters [30]. The primary established approach to enable reasoning in SLMs is the Finetune-CoT paradigm [16], which uses unstructured CoT explanations from the teacher as a supervision signal. This distillation approach has been demonstrated to be effective with multi-step mathematical reasoning [5]. However, relying on unstructured CoT often leads SLMs to learn spurious correlations rather than deep causal features, thereby reducing robustness on complex data [13, 29]. Consequently, recent research increasingly favors methods that enforce structural decomposition and explicit causality. These approaches include creating constrained distillation pipelines with the aim of removing ambiguous input context through advanced structural formats. For example, SocraticCoT provides an explicit structural format by guiding the model through reasoning using defined subquestion-solution pairs [29]. Alternatively, Mixed Distillation integrates the CoT with a formal verifiable Program-of-Thought reasoning to provide a less ambiguous supervision signal [13]. In the Text-to-SQL domain, structured KD has seen limited uptake. Previous studies, while validating the effectiveness of distillation in this area, have generally relied on less structured signals, such as schema-based finetuning to improve schema linking [11], the inclusion of enterprise-specific custom examples [10], and unstructured CoT reasoning transfer [27]. The proposed Struct-SQL framework systematically evaluates whether structure-based reasoning transfer, specifically QP-CoT, improves the distillation for Text-to-SQL.
5 Limitations
Struct-SQL significantly reduces syntactic errors through the use of structured reasoning; however, several constraints remain:
Token Overhead and Efficiency: Struct-SQL significantly reduces syntactic errors through structured reasoning; however, this approach introduces inference overhead. The generation of an intermediate query plan requires approximately 3.6 times more tokens than ReasonSQL (Table 3). This increase impacts latency and operational costs; nevertheless, these requirements may still be lower than those of teacher LLMs. It would be worthwhile to investigate more concise query plan templates to explore this trade-off.
Performance Bound: The student model’s capabilities are fundamentally capped by the teacher’s fidelity. Our distillation pipeline utilizes a strict filtering mechanism, training only on instances where the teacher model generates an execution-correct SQL query. Although this quality control is essential to prevent error propagation and provide a stable signal for small models [38], it inherently skews the training data toward problems that the teacher can solve, potentially limiting the student’s exposure to edge cases that the teacher itself cannot solve [25]. As a result, even with a competitive 60.42% EX on the BIRD test set, the performance remains bound by the teacher’s own limitations in handling highly complex multi-level joins and set operations, as discussed in Figure 4(a). To improve performance further, future work could explore using multiple teachers or human-annotated data to cover edge cases that a single source model may miss.
Benchmark Selection: A fixed reasoning template is used to minimize syntactic hallucinations through a constrained logical path; however, we recognize that it may limit flexibility across highly diverse database dialects. We intentionally prioritize evaluation on the BIRD benchmark, as it represents a more challenging dataset than SPIDER. As demonstrated in recent literature (e.g., Snowflake Arctic-Text2SQL-R1), state-of-the-art models frequently exceed 88% EX on the SPIDER-test, yet often struggle to surpass 70% on BIRD-dev [34]. This performance gap justifies our decision to focus exclusively on BIRD.
@@ 6 Conclusion and Future Research
This work presents Struct-SQL, a KD framework that transfers structured reasoning (QP-CoT) from a Teacher LLM to a smaller student model for the Text-to-SQL task. The central hypothesis is that a formal, structured reasoning signal provides a superior, less ambiguous teaching blueprint than unstructured reasoning. The experiments confirm the hypothesis, demonstrating that distilling structured reasoning is a better teaching method. Detailed error analysis indicates that these improvements are primarily due to a significant reduction in syntactic errors. An additional ablation study demonstrated generalization to a different base model. Given the demonstrated effectiveness of structured signals in the refinement of syntactic efficiency and logical reasoning, further research is warranted to validate the effect of this structured distillation approach on other complex reasoning tasks beyond Text-to-SQL. In summary, the proposed framework facilitates the deployment of high-performing, cost-effective and private models.
@@ 7 Acknowledgments
Appendix A Structured Reasoning Prompt Template (QP-CoT)
The QP-CoT prompt, as detailed in Listing 1, comprises a system instruction followed by two few-shot examples that demonstrate the query execution plan methodology. The prompt concludes with the target task; for the sake of brevity in this manuscript, the database schema definitions have been abbreviated.
Appendix B Unstructured Reasoning Prompt Template (CoT)
The ReasonSQL baseline (Listing 2) utilizes a 4-shot prompt to elicit free-form natural language reasoning. By requiring the model to ’think step-by-step’ without a formal execution plan, this baseline represents the standard unstructured distillation signal.
Appendix C Struct-SQL Output Examples
This section provides examples of the Struct-SQL model’s outputs across three levels of SQL complexity: Simple, Moderate, and Challenging. These examples demonstrate the model’s ability to generate a human-readable Query Plan that mirrors the logical execution steps of a database engine before producing the final SQLite SQL. Each entry includes the Question ID and Database ID for traceability. The following examples are selected based on the first successful generation encountered for each difficulty tier within the development dataset to ensure an unbiased representation of the model’s capabilities.